OpenStudy (anonymous):

find equations of the lines passing through (1,3) and have the following characteristics a) slope of -2/3 b) perpendicular to the line x+y=0 c) passing through the point (2,4) d) parallel to the x-axis

OpenStudy (zzr0ck3r):

what ones you need help with?

OpenStudy (anonymous):

all of them :/

OpenStudy (zzr0ck3r):

a) we know the equation for a line is y = mx+b where m is slope and b is y intercept. So 3 = (-2/3)*1 + b b = 3+2/3 b = 11/3 so a is y = (-2/3)x+11/3

OpenStudy (zzr0ck3r):

b)x+y = 0 so y = -x we know this slope is -1 and perpendicular means the negative recipicle of the slope so our new slope is 1 so 3 = 1 + b s0 b = 2 so our answer is y = x+2

OpenStudy (zzr0ck3r):

c) we now have two points (1,3) and (2,4) so our slope is (4-3)/(2-3) = -1 so choose a point to figure out b 3 = -1*1 + b b = 4 so y = -x+4

OpenStudy (zzr0ck3r):

d)the x axis is a slope of 0 so y = 0*x+b we use the point 3 = 1*0 + b b = 3 so y = 3 is the answer

OpenStudy (anonymous):

awesome, thanks!

OpenStudy (zzr0ck3r):

nonono you need to understand it, tell me what part you dont get?